# Effective synchronization amid noise-induced chaos

### v2

@@ III Convergent distributions
The synchronization scheme proposed in Sec. IV requires that $q_{\mathrm{A}}^{(k)}$ and $q_{\mathrm{B}}^{(k)}$ converge to the same distributions as $k\gg 1$. In this section, we demonstrate this convergence quantitatively via the decay of the KLD to zero for $\Lambda>0$.
@@
To this end, we simulate two oscillator ensembles, each consisting of $N=500$ phase samples. Their phase values are randomly and independently selected from different uniform distributions $\varphi^{(0)}_{{\mathrm{A}}n}\in[0,u)$ for the sender and $\varphi^{(0)}_{{\mathrm{B}}n}\in[0.5,0.5+u)$ for the receiver. By this, we study the extreme case of well-separated (non-overlapping) initial phase distributions. [^4] The true KLDs should have started from $\infty$ as the two initial distributions are nonoverlapping. Instead, in Figs. 3(a) and 4(a,b) they begin at a large finite value, $D^{(0)}\sim{-\log u}$. This is an artifact of the numerical KLD estimate (Eq. (10)). The exponential decay due to the mixing we report commences past that short transient, at which point the two distributions have the same support $\varphi\in[0,1)$, wherein indeed the KLD is a reliable measure. Other measures of similarity betweeen two distributions exist — notably the Wasserstein measure [36]. It is not singular for cases of non-overlapping supports, although it is only indirectly affected by the distributions’ peak heights; a central aspect to our analysis. Our choice of the KLD measure is motivated by our information-theoretic approach in terms of entropy. We subject the oscillators to Eq. (1) for given Lyapunov exponents $\Lambda$ and initial widths $u$. All $2N$ phases evolve under the same $\psi(\varphi)$ and uniformly-distributed waiting times $\beta^{(k)}\in[0,1)$. At each kick, we estimate $S^{(k)}_{\mathrm{A}}$, $S^{(k)}_{\mathrm{B}}$, $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$, and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$. [^5] The KLD is not symmetric. However, since the initial distributions are offset by exactly half of the phase circle and $\beta\in[0,1)$, we expect both $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$ and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$ to have the same statistics over many repetitions. Indeed their averages over realizations are equal in Fig. 4(a,b).
@@
For completeness, in SI Appendix, Sec. S5 we show how the synchronization for $\Lambda<0$ manifests in the entropy (see also Ref. [24]) and the KLD. For $\Lambda>0$, the two distributions $q_{\mathrm{A}}^{(k)}(\varphi)$ and $q_{\mathrm{B}}^{(k)}(\varphi)$ change erratically between consecutive kicks, e.g., as peaks are formed at phase values in the vicinity of the phase map’s extrema, and ‘smeared’ away from them (see Fig. 2 and SI Appendix, Fig. S4). Accordingly, in Fig. 3(b), we see that the entropy also changes chaotically [24]. At the same time, remarkably, both numerical KLDs decay to zero as seen in Fig. 3(a) despite the nonoverlapping supports of the initial distributions, indicating that the two ensembles eventually converge to the same distribution. [^6] The brief transient during $25\lesssim k\lesssim 65$, where the KLD seems to stay at a constant value is a result of the distributions being very narrow at these time steps (see Fig. 2), so they change widths together without affecting the KLD. We quantitatively explain this behavior in SI Appendix, Sec. S5. When the distribution widen afterwards, the KLD decreases to below its numerical error cutoff. Accordingly, the entropies of both distributions coincide at later times in Fig. 3(b). Note that a different realization of waiting times $\{\beta^{(k)}\}$ produces vastly different distributions (see SI Appendix, Fig. S5); what we report, therefore, is that the two agents’ distributions agree more closely with more kicks for every stochastic realization.
@@
We qualitatively understand this ensemble convergence as arising from independence of initial state. Consider the map from an initial phase $\varphi^{(0)}$ to the final phase $\varphi^{(k)}$ for a particular sequence of random and uniformly distributed waiting times $\{\beta^{(k)}\}$. This map comes from a statistically uniform sampling of the phase map $\psi(\varphi)$ and has an overall Lyapunov exponent for any $\varphi^{(0)}$ that approaches $k\Lambda$ as $k$ gets large. Since the phase circle is periodic and finite, nearby phases cannot be drifting apart as $\sim e^{k\Lambda}$ indefinitely. Instead, the non-monotonic phase map (Eq. (2)) “stretches and folds” distant phases akin to the Baker’s map, such that full mixing of $\varphi^{(0)}$s is possible [37]. [^7] This qualitative argument may suggest that full mixing is inevitable for such maps, though we know of neither a rigorous proof nor the necessary conditions for it to hold. Here, we have merely verified that the convergence in our examples was sufficient to produce closely-agreeing fiducial phases. As a result, the $\varphi^{(k)}$ for a given $\varphi^{(0)}$ may be arbitrarily close to the $\varphi^{\prime(k)}$ for other $\varphi^{\prime(0)}$’s residing anywhere throughout the phase circle. Then, any uncertainty in an observed $\varphi^{(k)}$ would erase all knowledge of $\varphi^{(0)}$. So, we expect that for any given uncertainty in $\varphi^{(k)}$ there exists some kick number $K_{\mathrm{m}}$ such that any dependence of $\varphi^{(k)}$ on $\varphi^{(0)}$ for $k>K_{\mathrm{m}}$ is undetectable [38]. This diminishing detectability would entail a KLD that approaches zero, implying that the two distributions differ only by this decaying uncertainty. We proceed to systematically show that the KLD in our examples indeed converges to zero.
@@ III.1 Convergence rate
Fig. 3 is a single realization of stochastic waiting times $\{\beta_{k}\}$. To find the typical time $K_{\mathrm{m}}$ that the agents should wait until their distributions become similar, we simulate a larger sampling of KLDs. To observe as many reliable KLD values as possible prior to their decay below the numerical-error cutoff of $\Delta D=0.05$, we picked the two initial distributions to be of width $u=10^{-4}$, so they are far apart and very narrow. We simulate $500$ realizations of $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$ and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$ dynamics, each repetition sampling the above initial distribution and picking $\{\beta^{(k)}\}$ anew. We plot, as an illustration, the average KLDs and their statistical uncertainty versus $k$ for $\Lambda=0.405$ in Fig. 4(a) and $\Lambda=0.121$ in Fig. 4(b).
@@
For all $\Lambda$’s, we observe a predominant regime of an exponential decay with $k$, shown with a dashed orange lines in Fig. 4(a,b). We anticipate a decay following the mixing argument of the preceding section. This exponential decay is typically cleaner for higher $\Lambda$’s (compare $k\in(10,45)$ in Fig. 4(a) with $k\in(70,250)$ in Fig. 4(b)), and lasts until we reach the numerical error cutoff for the KLD estimator, $\simeq 0.05$. Thus, for each $\Lambda$, we fit an exponential (as in Figs. 4(a,b)) to obtain a decay rate defined as $1/K_{\mathrm{m}}(\Lambda)$. We interpret these $K_{\mathrm{m}}$’s as the mixing-kick number. We verify for various $\Lambda$’s that it does not depend on the initial condition (by simulating $u=5\cdot 10^{-4}$ as well) or the sample size (by simulating $N=100$, too); $K_{\mathrm{m}}(\Lambda)$ remained identical within $3\%$ for all $\Lambda$’s, indicating that this convergence kick number is a robust quantity intrinsic to the dynamical system studied.
@@
In Fig. 4(c) we plot the mixing-kick number $K_{\mathrm{m}}$ versus $\Lambda$. For the cubic map, we numerically find that $K_{\mathrm{m}}$ strongly depends on $\Lambda$, varying across two orders of magnitude in a manner consistent with a power law, $K_{\mathrm{m}}(\Lambda)\sim\Lambda^{-1.491\pm 0.046}$. This implies that the more violent the phase map, [^8] By more violent we mean a larger gain parameter $A$ of the cubic map, Eq. (2). Then, when $\Lambda>0$, $\Lambda(A)$ is a monotonously increasing function for the cubic map; see also [12, Fig. 15.3]. the faster the mixing in the considered dynamics. We caution that the exponential fit for the decay of the KLD with low $\Lambda$ is not perfect (specifically, for $\Lambda\lesssim 0.2$), so their $K_{\mathrm{m}}$ should be taken with a grain of salt.
@@
We have no quantitative argument as to why $K_{\mathrm{m}}$ should diverge as $\Lambda\to 0^{+}$ as our findings suggest. Such a divergence is not obvious and should not apply generally, since the mixing property governing $K_{\mathrm{m}}$ is in principle different from the spreading property that governs $\Lambda$. We return to this point in Sec. V. In any case, the exact functional form of $K_{\mathrm{m}}(\Lambda)$ is not central to what follows; we will use the fact that $K_{\mathrm{m}}$ diverges as $\Lambda\to 0^{+}$ as a qualitative guide.
@@ IV Effective synchronization
In Sec. III, we found that after sufficiently many kicks, agents $\mathrm{A}$} and $\mathrm{B}$} effectively share $q^{(k)}(\varphi)$ in both the $\Lambda<0$ (synchronizing) and the $\Lambda>0$ (unsynchronizing) regimes. With $\Lambda<0$, the two agents agree at every moment (sufficiently-long after the most recent kick) on a specific phase to experimental accuracy and can use it to perform simultaneous actions or communicate. For $\Lambda>0$, while the resultant distributions do not asymptotically approach a sharp peak, the distributions of two agents become equivalent as numerically captured by the approach of the KLD among them to zero. The synchrony in the former and convergence in the latter remain with arbitrarily more kicks. Thus, any function of the sampled phases $\{\varphi^{(k)}_{1},\ldots,\varphi^{(k)}_{N}\}$ would be equal for both agents up to a statistical error. Below, we will consider a particular phase function we refer to as the “fiducial phase”, denoted $\varphi^{(k)}_{\mathrm{f}}=\varphi_{\mathrm{f}}[\varphi^{(k)}_{1},\ldots,\varphi^{(k)}_{N}$]. By phase function we mean that if $\{\varphi_{n}^{(k)}\}$ shift by $\Delta\varphi$, it must also shift by $\Delta\varphi$.
In order for agent $\mathrm{A}$} to act as though synchronized with agent $\mathrm{B}$}, it must be able to infer agent $\mathrm{B}$}’s fiducial phase $\varphi_{{\mathrm{f}},{\mathrm{B}}}$ at any given moment, given its phases $\{\varphi_{{\mathrm{A}}n}^{(k)}\}$. If agent $\mathrm{A}$}’s oscillator at that moment is $\varphi^{*}_{\mathrm{A}}$, its own fiducial phase $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}$ at that moment may be determined, since all phases advance at the same rate; see illustration in SI Appendix, Fig. S1. Specifically, $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}-\varphi^{*}_{\mathrm{A}}={\rm const}=\varphi^{(k)}_{{\mathrm{f}},{\mathrm{A}}}-\varphi^{(k)}_{\mathrm{A}}$; the same applies to agent $\mathrm{B}$}. Therefore, it remains to be shown that one may construct a function $\varphi^{(k)}_{{\mathrm{f}}}$ such that the estimated $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}=\varphi^{*}_{{\mathrm{f}},{\mathrm{B}}}$ to a good accuracy. We demonstrate below that this type of synchronization can indeed be achieved.
@@
To determine a peak position, one needs a smooth estimate of $q(\phi)$ from the sampled values $\{\varphi_{1},\ldots,\varphi_{N}\}$. We obtain an estimate denoted $Q(\varphi,\sigma)$ smoothed to a “bandwidth” $\sigma$ using the method of kernel-density estimation [39, 40, 41], detailed in Appendix A.2. The reliability of the $\varphi_{\mathrm{f}}$ obtained depends on $\sigma$. Large $\sigma$ produces a $Q(\varphi;\sigma)$ with a few broad peaks whose positions have high uncertainty, whereas small $\sigma$ gives numerous statistically insignificant peaks; see illustration in SI Appendix, Figs. S7.999Finding an optimal bandwidth $\sigma$, particularly for highly-multimodal distributions, is an open question in the field [42, 43, 44, 45].
Figure 5: (a) A density plot of the rescaled discrepancy, $\ell=\sqrt{N}|\Delta\varphi_{\mathrm{f}}|/e^{S}$ versus $S=(S_{\mathrm{A}}+S_{\mathrm{B}})/2$. The shade of blue represents the indicated values of the joint probability density function (PDF) to obtain $S$ and $\log_{10}\ell$, relative to its maximum. It was computed with kernel-density estimation from the $k=10^{4}$ data points shown in SI Appendix, Fig. S8(a), using a Gaussian kernel of width $0.4$ in $S$ and $0.2$ in $\log_{10}\ell$. The maximal value is $\mathrm{PDF}(-1.94,-0.0052)=0.135$, whose position is indicated by the black point. The puzzling lobe of probability at the upper right is addressed in SI Appendix, Sec. S8. (b) The marginal cumulative distribution function (CDF) of the discrepancy $\log_{10}\ell$, which was arranged into bins of size $0.25$. The dashed line is the error function of width unity; its close agreement with the data suggests that the discrepancy is predominantly normal-distributed with the expected scaling $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}/\sqrt{N}$. Inset: The marginal CDF of the entropy $S$, which was arranged into bins of size $0.5$. The dashed linear line implies, up to the low- and high-entropy outliers, that the entropy is exponentially distributed [24].
We observe that the multimodal distributions encountered for small positive $\Lambda$ have a dominant peak containing a substantial fraction of the $N$ samples (Fig. 2). In these narrowly-peaked distributions, the measured entropy $S$ (Eq. (6) and Eq. (9)) gives guidance about the appropriate $\sigma$: If all $\{\varphi_{n}\}$’s lie within an interval of width $w$, then $S=\ln w$ up to an additive constant [46]. Now, if instead a substantial fraction of them lie within $w$, the entropy remains comparable to $\ln w$. Thus, the measured entropy for each $q(\varphi)$ gives a characteristic length indicative of the dominant peak’s width. Thus, in what follows, we have simply used a bandwidth $\sigma$ equal to this $w$, i.e., $\sigma=e^{S}$, as the peak center will not be offset by more than its width. We explored other choices of $\sigma$ empirically to see their effect on the resulting uncertainties in $\varphi_{\mathrm{f}}$; see SI Appendix, Figs. S7 and S8. Accordingly, our fiducial phase of choice is
@@
This crude choice turns out to be sufficient for our purpose of demonstrating effective synchronization.

### v3

@@ III Convergent distributions
The synchronization scheme proposed in Sec. IV requires that $q_{\mathrm{A}}^{(k)}$ and $q_{\mathrm{B}}^{(k)}$ converge to the same distributions. In this section, we demonstrate this convergence quantitatively via the decay of the KLD to zero for $\Lambda>0$ after a finite number of kicks.
@@
To this end, we simulate two oscillator ensembles, each consisting of $N=500$ phase samples. Their phase values are randomly and independently selected from different uniform distributions $\varphi^{(0)}_{{\mathrm{A}}n}\in[0,u)$ for the sender and $\varphi^{(0)}_{{\mathrm{B}}n}\in[0.5,0.5+u)$ for the receiver. By this, we study the extreme case of well-separated (non-overlapping) initial phase distributions. [^5] The true KLDs should have started from $\infty$ as the two initial distributions are nonoverlapping. Instead, in Figs. 3(a) and 4(a,b) they begin at a large finite value, $D^{(0)}\sim{-\log u}$. This is an artifact of the numerical KLD estimate (Eq. (11)). The exponential decay due to the mixing we report commences past that short transient, at which point the two distributions have the same support $\varphi\in[0,1)$, wherein indeed the KLD is a reliable measure. We subject the oscillators to Eq. (1) for given Lyapunov exponents $\Lambda$ and initial widths $u$. All $2N$ phases evolve under the same $\psi(\varphi)$ and uniformly-distributed waiting times $\beta^{(k)}\in[0,1)$. At each kick, we estimate $S^{(k)}_{\mathrm{A}}$, $S^{(k)}_{\mathrm{B}}$, $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$, and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$. [^6] The KLD is not symmetric. However, since the initial distributions are offset by exactly half of the phase circle and $\beta\in[0,1)$, we expect both $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$ and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$ to have the same statistics over many repetitions. Indeed their averages over realizations are equal in Fig. 4(a,b).
@@
For completeness, in Appendix E we show how the synchronization for $\Lambda<0$ manifests in the entropy (see also Ref. [24]) and the KLD. For $\Lambda>0$, the two distributions $q_{\mathrm{A}}^{(k)}(\varphi)$ and $q_{\mathrm{B}}^{(k)}(\varphi)$ change erratically between consecutive kicks, e.g., as peaks are formed at phase values in the vicinity of the phase map’s extrema, and ‘smeared’ away from them (see Fig. 2 and Appendix C). Accordingly, in Fig. 3(b), we see that the entropy also changes chaotically [24]. At the same time, both numerical KLDs decay to zero as seen in Fig. 3(a) despite the distributions being different initially, indicating that the two ensembles eventually converge to the same $k$-dependent distributions. [^7] The brief transient during $25\lesssim k\lesssim 65$, where the KLD seems to stay at a constant value is a result of the distributions being very narrow at these time steps (see Fig. 2), so they change widths together without affecting the KLD. We quantitatively explain this behavior in Appendix E. When the distribution widen afterwards, the KLD decreases to below its numerical error cutoff. Accordingly, the entropies of both distributions coincide at later times in Fig. 3(b). Note that a different realization of waiting times $\{\beta^{(k)}\}$ produces vastly different distributions (see Appendix D), implying that common noise is really key to this effect.
@@
This ensemble convergence can be explained qualitatively as arising from independence of initial state. Consider the map from an initial phase $\varphi^{(0)}$ to the final phase $\varphi^{(k)}$ for a particular sequence of random and uniformly distributed waiting times $\{\beta^{(k)}\}$. This map comes from a statistically uniform sampling of the phase map $\psi(\varphi)$ and has an overall Lyapunov exponent for any $\varphi^{(0)}$ that approaches $k\Lambda$ as $k$ gets large. Since the phase circle is periodic and finite, nearby phases cannot be drifting apart as $\sim e^{k\Lambda}$ indefinitely. Instead, the non-monotonic phase map (Eq. (2)) “stretches and folds” phase samples, such that mixing of $\varphi^{(0)}$s is possible. As a result, the $\varphi^{(k)}$ for a given $\varphi^{(0)}$ may be arbitrarily close to the $\varphi^{\prime(k)}$ for other $\varphi^{\prime(0)}$’s residing anywhere throughout the phase circle. Then, any uncertainty in an observed $\varphi^{(k)}$ would erase all knowledge of $\varphi^{(0)}$. The asymptotic convergence of phase distributions under stochastic maps has been rigorously proven for general classes of stochastic systems [29, 30, 31, 32, 40], based on spreading-of-phases arguments akin to the above.
To demonstrate applicability, however, we must go a step further and show that for any given uncertainty in $\varphi^{(k)}$ there exists some well-defined and finite kick number $K_{\mathrm{m}}$ such that any dependence of $\varphi^{(k)}$ on $\varphi^{(0)}$ for $k>K_{\mathrm{m}}$ is undetectable [41]. This diminishing detectability would entail a KLD that approaches zero, implying that the two distributions differ only by this decaying uncertainty. We thus proceed to confirm numerically that the KLD in our examples indeed converges to zero after a systematically predictable mixing number of kicks.
@@ III.1 Convergence rate
Fig. 3 is a single realization of stochastic waiting times $\{\beta^{(k})\}$. To find a typical time $K_{\mathrm{m}}$ that the agents should wait until their distributions become similar, we simulate a larger sampling of KLDs. To observe as many reliable KLD values as possible prior to their decay below the numerical-error cutoff of $\Delta D=0.05$, we picked the two initial distributions to be of width $u=10^{-4}$, so they are far apart and very narrow. We simulate $500$ realizations of $D^{(k)}({\mathrm{A}}\|{\mathrm{B}})$ and $D^{(k)}({\mathrm{B}}\|{\mathrm{A}})$ dynamics, each repetition sampling the above initial distribution and picking $\{\beta^{(k)}\}$ anew. We plot, as an illustration, the average KLDs and their statistical uncertainty versus $k$ for $\Lambda=0.405$ in Fig. 4(a) and $\Lambda=0.121$ in Fig. 4(b).
@@
For all $\Lambda$’s, we observe a predominant regime of an exponential decay with $k$, shown with a dashed orange lines in Fig. 4(a,b). We anticipate a decay following the mixing argument of the preceding section. This exponential decay is typically cleaner for higher $\Lambda$’s (compare $k\in(10,45)$ in Fig. 4(a) with $k\in(70,250)$ in Fig. 4(b)), and lasts until we reach the numerical error cutoff for the KLD estimator, $\simeq 0.05$. Thus, for each $\Lambda$, we fit an exponential (as in Figs. 4(a,b)) to obtain a decay rate defined as $1/K_{\mathrm{m}}(\Lambda)$. We interpret these $K_{\mathrm{m}}$’s as the mixing-kick number. We verify for all $\Lambda$’s that it does not depend on the initial condition (by simulating $u=5\cdot 10^{-4}$ as well) or the sample size (by simulating $N=100$, too); $K_{\mathrm{m}}(\Lambda)$ remained identical within $3\%$ for all $\Lambda$’s, indicating that this convergence kick number is a robust quantity intrinsic to the dynamical system studied.
@@
In Fig. 4(c) we plot the mixing-kick number $K_{\mathrm{m}}$ versus $\Lambda$. For the cubic map, we numerically find that $K_{\mathrm{m}}$ strongly depends on $\Lambda$, varying across two orders of magnitude in a manner consistent with a power law, $K_{\mathrm{m}}(\Lambda)\sim\Lambda^{-1.491\pm 0.046}$. This implies that the more violent the phase map, [^8] By more violent we mean a larger gain parameter $A$ of the cubic map, Eq. (2). For $\Lambda>0$, $\Lambda(A)$ is a monotonously increasing function for the cubic map; see, e.g. Fig. 15.3 in Ref. [12]. the faster the mixing in the considered dynamics. We caution that the exponential fit for the decay of the KLD with low $\Lambda$ is not perfect (specifically, for $\Lambda\lesssim 0.2$), so their $K_{\mathrm{m}}$ should be taken with a grain of salt.
@@
We have no quantitative argument as to why $K_{\mathrm{m}}$ should diverge as $\Lambda\to 0^{+}$ as our findings suggest. Generally, the mixing property governing $K_{\mathrm{m}}$ is in principle different from the spreading property that governs $\Lambda$. Qualitatively, the spreading of phases seems to be the only effect enabling mixing, so the absence of spreading ($\Lambda\to 0^{+}$) should imply the absence of mixing ($1/K_{\mathrm{m}}\to 0$). From a different perspective, we point out in Sec. V.2 that $\Lambda=0$ marks the transition between a random point attractor and an extended random attractor [40]. In any case, the exact functional form of $K_{\mathrm{m}}(\Lambda)$ is not central to what follows; we will use the fact that $K_{\mathrm{m}}$ diverges as $\Lambda\to 0^{+}$ and that it is well-defined and finite otherwise as a qualitative guide.
@@ IV Effective synchronization
In Sec. III, we found that after finitely many kicks, agents $\mathrm{A}$} and $\mathrm{B}$} effectively share $q^{(k)}(\varphi)$ in both the $\Lambda<0$ (synchronizing) and the $\Lambda>0$ (unsynchronizing) regimes. With $\Lambda<0$, the two agents agree at every moment (sufficiently-long after the most recent kick) on a specific phase to experimental accuracy and can use it to perform simultaneous actions or communicate. For $\Lambda>0$, while the resultant distributions do not asymptotically approach a sharp peak, the distributions of two agents become equivalent as numerically captured by the approach of the KLD among them to zero. The synchrony in the former and convergence in the latter remain with arbitrarily more kicks. Thus, any function of the sampled phases $\{\varphi^{(k)}_{1},\ldots,\varphi^{(k)}_{N}\}$ would be equal for both agents up to a statistical error. Below, we will consider a particular phase function we refer to as the “fiducial phase”, denoted $\varphi^{(k)}_{\mathrm{f}}=\varphi_{\mathrm{f}}[\varphi^{(k)}_{1},\ldots,\varphi^{(k)}_{N}$]. By phase function we mean that if $\{\varphi_{n}^{(k)}\}$ shift by $\Delta\varphi$, it must also shift by $\Delta\varphi$.
In order for agent $\mathrm{A}$} to act as though synchronized with agent $\mathrm{B}$}, it must be able to infer agent $\mathrm{B}$}’s fiducial phase $\varphi_{{\mathrm{f}},{\mathrm{B}}}$ at any given moment, given its phases $\{\varphi_{{\mathrm{A}}n}^{(k)}\}$. If agent $\mathrm{A}$}’s oscillator at that moment is $\varphi^{*}_{\mathrm{A}}$, its own fiducial phase $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}$ at that moment may be determined, since all phases advance at the same rate; see Appendix F. Specifically, $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}-\varphi^{*}_{\mathrm{A}}={\rm const}=\varphi^{(k)}_{{\mathrm{f}},{\mathrm{A}}}-\varphi^{(k)}_{\mathrm{A}}$; the same applies to agent $\mathrm{B}$}. Therefore, it remains to be shown that one may construct a function $\varphi^{(k)}_{{\mathrm{f}}}$ such that the estimated $\varphi^{*}_{{\mathrm{f}},{\mathrm{A}}}=\varphi^{*}_{{\mathrm{f}},{\mathrm{B}}}$ to a good accuracy. We demonstrate below that this type of synchronization can indeed be achieved.
@@
To determine a peak position, one needs a smooth estimate of $q(\varphi)$ from the sampled values $\{\varphi_{1},\ldots,\varphi_{N}\}$. We obtain an estimate denoted $Q(\varphi,\sigma)$ smoothed to a “bandwidth” $\sigma$ using the method of kernel-density estimation [42, 43, 44], detailed in Appendix G. The reliability of the $\varphi_{\mathrm{f}}$ obtained depends on $\sigma$. Large $\sigma$ produces a $Q(\varphi;\sigma)$ with a few broad peaks whose positions have high uncertainty, whereas small $\sigma$ gives numerous statistically insignificant peaks; see Appendix H. [^9] Finding an optimal bandwidth $\sigma$, particularly for highly-multimodal distributions, is an open question in the field [45, 46, 47, 48].
We observe that the multimodal distributions encountered for small positive $\Lambda$ have a dominant peak containing a substantial fraction of the $N$ samples (Fig. 2). In these narrowly-peaked distributions, the measured entropy $S$ (Eq. (6) and Eq. (10)) gives guidance about the appropriate $\sigma$: If all $\{\varphi_{n}\}$’s lie within an interval of width $w$, then $S=\ln w$ up to an additive constant [49]. Now, if instead a substantial fraction of them lie within $w$, the entropy remains comparable to $\ln w$. Thus, the measured entropy for each $q(\varphi)$ gives a characteristic length indicative of the dominant peak’s width. Thus, in what follows, we have simply used a bandwidth $\sigma$ equal to this $w$, i.e., $\sigma=e^{S}$, as the peak center will not be offset by more than its width. We explored other choices of $\sigma$ empirically to see their effect on the resulting uncertainties in $\varphi_{\mathrm{f}}$; see Appendixes H and I. Accordingly, our fiducial phase of choice is
@@
Even with this crude choice, we are able to demonstrate a clear effective synchronization.